# Quantum Computing for Energy Correlators

### v1

@@ IV.1.2 $5\times 5$ lattice
Figure 14: Energy conservation test on a $5\times 5$ honeycomb lattice with $ag^{2}=1$ and $j_{\rm max}=\frac{1}{2}$: (a) energy of the central plaquette marked in orange in Fig. 12 plus the integrated total energy flux going out of the orange plaquette; (b) energies of the plaquettes marked in orange and blue in Fig. 12 plus the integrated total energy flux going out of the blue plaquettes.
@@
We first study the time dependence of the expectation value of the energy flux operator as defined in Eq. (54). We first look at the energy flux from the center into blue plaquettes similar to what we have already studied for the $3\times 3$ lattice. In Fig. 13a, we notice that the boundary effects are significantly reduced compared to the $3\times 3$ case and the rotation symmetry is more respected. In Fig. 13b, we also study the energy flux into the green plaquettes, which we can consider for the first time on the $5\times 5$ lattice. As the energy flux originating from the center ends up on green plaquettes only by going through blue plaquettes first, it is sufficient to look at the energy flux from blue to green. The energy flux along $(3,2)\to(4,2)$ and $(2,3)\to(2,4)$ exactly agree, which is a result of the reflection symmetry of the lattice system along the axis connecting $(0,0)$ and $(4,4)$. The energy flux from $(3,1)$ to $(4,0)$ agrees with the other two approximately for the time period studied here, which indicates the rotational symmetry is approximately respected. The agreement is significantly improved between $t=13$ and $t=20$, compared with the $3\times 3$ case in Fig. 8. We attribute this improvement to the larger lattice with smaller boundary effects.
@@
We further test energy conservation on the $5\times 5$ lattice as we did for $3\times 3$ in Fig. 9. We perform two tests and present the results in Fig. 14. In the first one, we calculate the energy on the central plaquette marked in orange at $(2,2)$ as a function of time and the integrated total energy flux going from the orange plaquette to the blue plaquettes. In the second test, we compare the energies summed over the orange and blue plaquettes and the integrated total energy flux going from the blue to the green plaquettes. In both tests, energy conservation is well maintained. The total energy of the orange and blue plaquettes at $t=0$ is slightly below the energy of the orange plaquette at $t=0$, since the energy density of the ground state is slightly below zero in our energy reference convention of the Hamiltonian in Eq. (III.3).
@@
Figure 15: Two-point energy correlators as functions of the coupling at two different relative angles: $\theta_{12}=\pi$ (black dots) and $\theta_{12}=\frac{2\pi}{3}$ (red crosses), compared with the one-point energy correlator squared (blue stars) on a $5\times 5$ honeycomb lattice with $j_{\rm max}=\frac{1}{2}$ and a local source at $(2,2)$. All energy correlators are calculated from the green sites solely, as marked in Fig. 12. Since the detectors are separated from the local source by one lattice site, the lattice artifact of vanishing $\langle\mathcal{E}_{1}\rangle$ around $ag^{2}=1$ as seen in Figs. 10 and 11 disappears here.
@@
Finally, we calculate two-point energy correlators on the $5\times 5$ lattice. In Fig. 15, we show the energy-energy correlators with $\theta_{12}=\pi$ and $32\pi/$ 3 as functions of the coupling, compared with the one-point energy correlator squared. To compute energy correlators with such angles, we use rotational symmetry and fix specific plaquette sites with such angular separation, rather than compute all the pairs with the same angular separation. In particular, we choose $\vec{n}_{1}:(3,1)\to(4,0)$ and $\vec{n}_{2}:(1,3)\to(0,4)$ or $\vec{n}_{2}:(2,3)\to(2,4)$, i.e. compute the correlation between $(4,0)$ and $(0,4)$ or $(4,0)$ and $(2,4)$. We also compute the one-point correlator along $\vec{n}_{1}$. We integrate from $t=0$ to $t=0.1$. Approximate factorization is observed for both angles at all the couplings studied here, indicating no jets on our current small and highly truncated lattice. The lattice artifact, i.e., the vanishing of the one-point energy correlator around $ag^{2}\approx 1$ that we observed for $3\times 3$ in Figs. 10 and 11 disappears for $5\times 5$.
(a) Results for $ag^{2}=0.6$ with the two detectors at $(2,0)$ and $(0,2)$.
(b) Results for $ag^{2}=0.6$ with the two detectors at $(2,0)$ and $(1,2)$.
@@
Figure 16: Results from IBM Aer emulator running all 36 quantum circuits for computing the real part of the energy-energy correlator for different detector positions on a $3\times 3$ lattice with $j_{\rm max}=\frac{1}{2}$ at two couplings $ag^{2}=0.6$ (top row) and $ag^{2}=1$ (bottom row). The local source is located at plaquette $(1,1)$. Blue points indicate the quantum circuit results obtained from the emulator, the solid black lines label the exact numerical results, and the green dashed lines represent the exact quantum circuit results (including the Trotterization errors in both adiabatic evolution and real-time evolution). The energy flux operator is given by Eq. (III.3), which is implemented as described in Sec. III.4.
@@ IV.2 Quantum results from Aer emulator
We place the two detectors at different locations: for $ag^{2}=1$, the first one at $(i,j)=(0,2)$ and the second one at $(i,j)=(2,1)$ or $(2,0)$; for $ag^{2}=0.6$, the first one at $(i,j)=(2,0)$ and the second one at $(i,j)=(1,2)$ or $(0,2)$. Since our main purpose here is to test the quantum algorithm for calculating the real-time Wightman correlators, rather than obtaining physical results, we will also allow time ranges where the two detectors are no longer spacelike separated. As a result, the energy correlator may become complex. We will only test the quantum algorithms for computing real part of the energy correlator, though the algorithm to calculate the imaginary part was also described in Fig. 4. In particular, we consider $t_{1}\in[0,2]$ for the first detector with $t_{2}=1$ fixed for the second detector. In the Trotterization for the real-time evolution, we set $\Delta t=0.025$. As mentioned in Sec. III.4, we need to run 36 different quantum circuits to calculate the two-point real-time correlator.
Fig. 16 shows the obtained results for the energy-energy correlator from the Aer emulator using $2\times 10^{6}$ shots. The top panels show the results for $ag^{2}=0.6$ while the lower panels for $ag^{2}=1$. We observe that the results from the quantum emulator (blue squares) give results compatible within two sigmas with the exact numerical results (solid black lines) and the noiseless simulation results (dashed green lines). The errorbars associated with the emulator results are statistical, which are estimated by using binomial distribution and standard uncertainty propagation. The difference between the exact numerical results and the noiseless simulation results originates from the Trotterization in the adiabatic and real-time evolution. We note that in the $ag^{2}=0.6$ case, the real part of the real-time correlator has a bigger value and thus the statistical uncertainty of using $2\times 10^{6}$ shots is smaller than that in the $ag^{2}=1$ case using the same number of shots. This indicates that one needs to use more shots for the $ag^{2}=1$ case in order to achieve the same accuracy as in the $ag^{2}=0.6$ case.
@@ Acknowledgements.
We thank Adam Levine, Ian Moult, Natalie Klco, Martin Savage, Matthew Walters, and Nikita Zemlevskiy for useful discussions. K.L. was supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics from DE-SC0011090. F.T. and X.Y. were supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics, InQubator for Quantum Simulation (IQuS) (https://iqus.uw.edu) under Award Number DOE (NP) Award DE-SC0020970 via the program on Quantum Horizons: QIS Research and Innovation for Nuclear Science. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a Department of Energy Office of Science User Facility using NERSC award NP-ERCAP0027114.
@@ Appendix A Bare stress-energy tensor on the honeycomb lattice
On the honeycomb lattice, we can identify electric field and plaquette operator as
@@
Using this identification naively, we can obtain the bare lattice version of the energy flux operator given in Eq. (36). Then the energy flux operator with spatial components $x$ and $y$ are given as
@@
$\qquad\qquad-E_{x}^{a}{\rm Tr}(T^{a}U_{\varhexagon}^{\dagger})+{\rm Tr}(T^{a}U_{\varhexagon}^{\dagger})E_{x}^{a}\big{)}\,,$ (58)
@@
Second, $\hat{x}$ and $\hat{y}$ are not the natural directions of the honeycomb lattice. As shown in Fig. 6, the three natural directions for electric fields on the links are denoted as $\hat{e}_{i}$ for $i=1,2,3$. We now try to construct energy flux operator using the directions more natural to the honeycomb lattice.
We first note that the above expression of $T^{\rm bare}_{0x}$ contains only $E_{y}^{a}$ while that of $T^{\rm bare}_{0y}$ contains only $E_{x}^{a}$, indicating that the direction of the energy flux operator is perpendicular to that of the electric field, as in the Poynting vector. This observation informs us how to generalize Eq. (A) for the honeycomb lattice: the electric field flowing on a given link describes the energy flux across that link.
For a plaquette located at position $(i,j)$ as in Fig. 7, the six possible energy flux directions $(i^{\prime},j^{\prime})$ are given as $(i^{\prime},j^{\prime})\in\{(i-1,j+1),(i,j+1),(i+1,j),(i+1,j-1),(i,j-1),(i-1,j)\}$, corresponding directly to the energy flux across the six links labeled 1 through 6, respectively. As discussed around Eq. (37), we can label associated energy flux operator between $(i,j)$ and $(i^{\prime},j^{\prime})$ as $T^{\rm bare}_{0,(i,j)\to(i^{\prime},j^{\prime})}$. For example, if we consider electric fields pointing from red vertex along the link 1 and link 6, we find these electric fields are related to the energy fluxes between $(i,j)\to(i-1,j+1)$ and $(i,j)\to(i-1,j)$, respectively. As shown in Fig. 6, these two directions from the red vertex are denoted as $\hat{e}_{1}$ and $\hat{e}_{3}$ respectively, and thus the electric fields along them starting from the red vertex are denoted as $E_{R1}^{a}$ and $E_{R3}^{a}$, respectively. As our plaquette operator $U_{\varhexagon}$ is defined with the convention that traces over six gauge links in counter-clockwise direction, it is useful to consider both electric fields pointing from the red vertex along the link 1 in $\hat{e}_{1}$ direction (counter-clockwise direction) and along the link 6 in $\hat{e}_{3}$ direction (clockwise direction).
Shifting the $E_{x}^{a}$ or $E_{y}^{a}$ from Eq. (A) to $E_{R1}^{a}$ and $E_{R3}^{a}$, which respectively indicates the electric field pointing along the link 1 and 6 from the red vertex, the relevant matrix elements of the energy flux operator in the physical Hilbert space is given by for link 1 as
@@
where repeated indices are summed over. The internal and external link states are specified as in Fig. 7. The big parenthesis with six numbers is the Wigner-$3j$ symbol. Similarly, the expression for the electric field along link 6 from the red vertex is given as
@@
Finally, we note that we also need terms involving $U_{\varhexagon}^{\dagger}$ as indicated in Eq. (A). We note that in the plaquette at $(i,j)$ shown in Fig. 7, the Wilson line on link 1 inside $U_{\varhexagon}$ starts on the red dot, i.e., it is $U_{1}$. The $U_{\varhexagon}^{\dagger}$ as needed in Eq. (A) is obtained from the plaquette at $(i-1,j+1)$ since the Wilson line on link 1 there ends on the red dot, i.e., it is $U_{1}^{\dagger}$. Similar argument applies to link 6. Putting all these together, we find the energy flux from plaquette $(i,j)$ to plaquette $(i-1,j+1)$ or $(i-1,j)$ is given by per plaquette area in lattice units)

### v2

@@ IV.1.2 $5\times 5$ lattice
Figure 14: Energy conservation test on a $5\times 5$ honeycomb lattice with $ag^{2}=1$ and $j_{\rm max}=\frac{1}{2}$: (a) energy of the central plaquette marked in orange in Fig. 12 plus the integrated total energy flux going out of the orange plaquette; (b) energies of the plaquettes marked in orange and blue in Fig. 12 plus the integrated total energy flux going from the blue to the green plaquettes.
@@
We first study the time dependence of the expectation value of the energy flux operator as defined in Eq. (54). We first look at the energy flux from the center into blue plaquettes similar to what we have already studied for the $3\times 3$ lattice. In Fig. 13a, we notice that the boundary effects are significantly reduced compared to the $3\times 3$ case and the rotation symmetry is more respected. In Fig. 13b, we also study the energy flux into the green plaquettes, which we can consider for the first time on the $5\times 5$ lattice. As the energy flux originating from the center ends up on green plaquettes only by going through blue plaquettes first, it is sufficient to look at the energy flux from blue to green plaquettes. The energy fluxes along $(3,2)\to(4,2)$ and $(2,3)\to(2,4)$ agree exactly, which is a result of the reflection symmetry of the lattice system along the axis connecting $(0,0)$ and $(4,4)$. The energy flux from $(3,1)$ to $(4,0)$ agrees with the other two approximately for the time period studied here, which indicates the rotational symmetry is approximately respected. The agreement is significantly improved between $t=13$ and $t=20$, compared with the $3\times 3$ case in Fig. 8. We attribute this improvement to the larger lattice with smaller boundary effects.
@@
We further test energy conservation on the $5\times 5$ lattice as we did for $3\times 3$ in Fig. 9. We perform two tests and present the results in Fig. 14. In the first one, we calculate the energy on the central plaquette marked in orange at $(2,2)$ as a function of time and the integrated total energy flux going from the orange plaquette to the blue plaquettes. In the second test, we compare the energies summed over the orange and blue plaquettes and the integrated total energy flux going from the blue to the green plaquettes. In both tests, energy conservation is well maintained. The total energy of the orange and blue plaquettes at $t=0$ is slightly below the energy of the orange plaquette at $t=0$, since the energy density of the ground state is slightly below zero in our convention of the energy reference of the Hamiltonian in Eq. (III.3).
@@
Figure 15: Two-point energy correlators as functions of the coupling at two different relative angles: $\theta_{12}=\pi$ (black dots) and $\theta_{12}=\frac{2\pi}{3}$ (red crosses), compared with the one-point energy correlator squared (blue stars) on a $5\times 5$ honeycomb lattice with $j_{\rm max}=\frac{1}{2}$ and a local source at $(2,2)$. All energy correlators are calculated from the fluxes going from the blue to green sites, as marked in Fig. 12. Since the detectors are separated from the local source by one lattice site, the lattice artifact of vanishing $\langle\mathcal{E}_{1}\rangle$ around $ag^{2}=1$ as seen in Figs. 10 and 11 disappears here.
@@
Finally, we calculate two-point energy correlators on the $5\times 5$ lattice. In Fig. 15, we show the energy-energy correlators with $\theta_{12}=\pi$ and $32\pi/$ 3 as functions of the coupling, compared with the square of the one-point energy correlator. To compute energy correlators with such angles, we use rotational symmetry and fix specific plaquette sites with such angular separation, rather than compute all the pairs of sites with the same angular separation. In particular, we choose $\vec{n}_{1}:(3,1)\to(4,0)$ and $\vec{n}_{2}:(1,3)\to(0,4)$ or $\vec{n}_{2}:(2,3)\to(2,4)$, i.e., compute the correlation between $(4,0)$ and $(0,4)$ or $(4,0)$ and $(2,4)$. We also compute the one-point correlator along $\vec{n}_{1}$. We integrate from $t=0$ to $t=0.1$. Approximate factorization is observed for both angles at all the couplings studied here, indicating no jets on our current small and highly truncated lattice. The lattice artifact, i.e., the vanishing of the one-point energy correlator around $ag^{2}\approx 1$ that we observed on the $3\times 3$ lattice in Figs. 10 and 11 disappears on the $5\times 5$.
(a) Results for $ag^{2}=0.6$ with the two detectors at $(0,2)$ and $(2,0)$.
(b) Results for $ag^{2}=0.6$ with the two detectors at $(0,2)$ and $(2,1)$.
@@
Figure 16: Results from IBM Aer emulator running all 36 quantum circuits for computing the real part of the energy-energy correlator for different detector positions on a $3\times 3$ lattice with $j_{\rm max}=\frac{1}{2}$ at two couplings $ag^{2}=0.6$ (top row) and $ag^{2}=1$ (bottom row). The local source is located at plaquette $(1,1)$. Blue points indicate the quantum circuit results obtained from the emulator, the black solid lines label the exact numerical results, and the green dashed lines represent the exact quantum circuit results (including the Trotterization errors in both adiabatic evolution and real-time evolution). The energy flux operator is given by Eq. (III.3), which is implemented as described in Sec. III.4.
@@ IV.2 Quantum results from Aer emulator
We place the two detectors at different locations: the first one at $(i,j)=(0,2)$ and the second one at $(i,j)=(2,1)$ or $(2,0)$. Since our main purpose here is to test the quantum algorithm for calculating the real-time Wightman correlators, rather than obtaining physical results, we will also allow time ranges where the two detectors are no longer spacelike separated. As a result, the energy correlator may become complex. We will only test the quantum algorithms that compute the real part of the energy correlator, though the algorithm to calculate the imaginary part was also described in Fig. 4. In particular, we consider $t_{1}\in[0,2]$ for the first detector with $t_{2}=1$ fixed for the second detector. In the Trotterization of the real-time evolution, we set $\Delta t=0.025$. As mentioned in Sec. III.4, we need to run 36 different quantum circuits to calculate the two-point real-time correlator.
Figure 16 shows the results of the energy-energy correlators obtained from the Aer emulator using $2\times 10^{6}$ shots. The top panels show the results for $ag^{2}=0.6$ while the lower panels for $ag^{2}=1$. We observe that the results from the quantum emulator (blue squares) give results compatible within two sigmas with the exact numerical results (black solid lines) and the noiseless simulation results (green dashed lines). The errorbars associated with the emulator results are statistical, which are estimated by using binomial distribution and standard uncertainty propagation. The difference between the exact numerical results and the noiseless simulation results originates from the Trotterization in the adiabatic and real-time evolution. We note that in the $ag^{2}=0.6$ case, the real part of the real-time correlator has a bigger value and thus the statistical uncertainty of using $2\times 10^{6}$ shots is smaller than that in the $ag^{2}=1$ case using the same number of shots. This indicates that one needs to use more shots for the $ag^{2}=1$ case in order to achieve the same accuracy as in the $ag^{2}=0.6$ case.
@@ Acknowledgements.
We thank Adam Levine, Ian Moult, Natalie Klco, Martin Savage, Matthew Walters, Neill Warrington, and Nikita Zemlevskiy for useful discussions. K.L. was supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics from DE-SC0011090. F.T. and X.Y. were supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics, InQubator for Quantum Simulation (IQuS) (https://iqus.uw.edu) under Award Number DOE (NP) Award DE-SC0020970 via the program on Quantum Horizons: QIS Research and Innovation for Nuclear Science. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a Department of Energy Office of Science User Facility using NERSC award NP-ERCAP0027114.
@@ Appendix A Bare stress-energy tensor on the honeycomb lattice
On the honeycomb lattice, we can identify the electric field and the plaquette operator as
@@
Using this identification naively, we can obtain the bare lattice version of the energy flux operator defined in Eq. (36). Its $x$ and $y$ spatial components are given as
@@
$\qquad\qquad-E_{x}^{a}{\rm Tr}(T^{a}U_{\varhexagon}^{\dagger})-{\rm Tr}(T^{a}U_{\varhexagon}^{\dagger})E_{x}^{a}\big{)}\,,$ (58)
@@
Second, $\hat{x}$ and $\hat{y}$ are not the natural directions of the honeycomb lattice. As shown in Fig. 6, the three natural directions for electric fields on the links are denoted as $\hat{e}_{i}$ for $i=1,2,3$. We now try to construct energy flux operators using the directions more natural to the honeycomb lattice.
We note that the above expression of $T^{\rm bare}_{0x}$ contains only $E_{y}^{a}$ while that of $T^{\rm bare}_{0y}$ contains only $E_{x}^{a}$, indicating that the direction of the energy flux operator is perpendicular to that of the electric field, as in the Poynting vector. This observation informs us how to generalize Eq. (A) for the honeycomb lattice: the expression containing the electric field on a given link describes the energy flux across that link.
For a plaquette located at position $(i,j)$ as in Fig. 7, the six possible energy flux directions $(i^{\prime},j^{\prime})$ are given as $(i^{\prime},j^{\prime})\in\{(i-1,j+1),(i,j+1),(i+1,j),(i+1,j-1),(i,j-1),(i-1,j)\}$, corresponding directly to the energy flux across the six links labeled 1 through 6, respectively. As discussed around Eq. (37), we can label associated energy flux operator between $(i,j)$ and $(i^{\prime},j^{\prime})$ as $T^{\rm bare}_{0,(i,j)\to(i^{\prime},j^{\prime})}$. For example, if we consider electric fields pointing from the red vertex along the link 1 and link 6, we find these electric fields are related to the energy fluxes between $(i,j)\to(i-1,j+1)$ and $(i,j)\to(i-1,j)$, respectively. As shown in Fig. 6, these two directions on link 1 and 6 are denoted as $\hat{e}_{1}$ and $\hat{e}_{3}$, and thus the electric fields along them starting from the red vertex are denoted as $E_{R1}^{a}$ and $E_{R3}^{a}$, respectively. As our plaquette operator $U_{\varhexagon}$ is defined with the convention that traces over six gauge links in counter-clockwise direction, it is useful to consider both electric fields pointing from the red vertex along the link 1 in $\hat{e}_{1}$ direction (counter-clockwise direction) and along the link 6 in $\hat{e}_{3}$ direction (clockwise direction).
Shifting the $E_{x}^{a}$ or $E_{y}^{a}$ from Eq. (A) to $E_{R1}^{a}$ and $E_{R3}^{a}$, which respectively indicates the electric field pointing along the link 1 and 6 from the red vertex, the relevant matrix elements of the energy flux operator for link 1 in the physical Hilbert space is given by
@@
where repeated indices are summed over. The internal and external link states are specified as in Fig. 7. The big parenthesis with six numbers is the Wigner-$3j$ symbol. Similarly, the expression of the energy flux associated with the electric field along link 6 from the red vertex is given by
@@
Finally, we note that we also need terms involving $U_{\varhexagon}^{\dagger}$ as indicated in Eq. (A). In the plaquette at $(i,j)$ shown in Fig. 7, the Wilson line on link 1 inside $U_{\varhexagon}$ starts on the red dot, i.e., it is $U_{1}$. The $U_{\varhexagon}^{\dagger}$ as needed in Eq. (A) is obtained from the plaquette at $(i-1,j+1)$ since the Wilson line on link 1 there ends on the red dot, i.e., it is $U_{1}^{\dagger}$. A similar argument applies to link 6. Putting all these together, we find the energy flux from plaquette $(i,j)$ to plaquette $(i-1,j+1)$ or $(i-1,j)$ is given by per plaquette area in lattice units